### v2

# MineLand: Simulating Large-Scale Multi-Agent Interactions with Limited Multimodal Senses and Physical Needs

Abstract
While Vision-Language Models (VLMs) hold promise for tasks requiring extensive collaboration, traditional multi-agent simulators have facilitated rich explorations of an interactive artificial society that reflects collective behavior. However, these existing simulators face significant limitations. Firstly, they struggle with handling large numbers of agents due to high resource demands. Secondly, they often assume agents possess perfect information and limitless capabilities, hindering the ecological validity of simulated social interactions. To bridge this gap, we propose a multi-agent Minecraft simulator, MineLand, that bridges this gap by introducing three key features: large-scale scalability, limited multimodal senses, and physical needs. Our simulator supports 64 or more agents. Agents have limited visual, auditory, and environmental awareness, forcing them to actively communicate and collaborate to fulfill physical needs like food and resources. Additionally, we further introduce an AI agent framework, Alex, inspired by multitasking theory, enabling agents to handle intricate coordination and scheduling. Our experiments demonstrate that the simulator, the corresponding benchmark, and the AI agent framework contribute to more ecological and nuanced collective behavior. The source code of MineLand and Alex is openly available at https://github.com/cocacola-lab/MineLand.
Figure 1: A panoramic view of one scene in MineLand, consisting of multiple AI agents. Subfigure 3&6 show interactions demonstrating cooperation and competition among several agents. Subfigure 5&2&4 showcases the scenarios where the limited senses, physical needs, and multitasking mechanism reflect. In Subfigure 1, an agent is performing a creative task named Exploration. Two agents in the left cave of Subfigure 5 cooperate to finish a programmatic mining task, while agents in Subfigure 3 are carrying out building construction, which is a hybrid task.
1 Introduction
Multi-agent simulators have facilitated rich explorations of an interactive artificial society that reflects collective behavior. From sandbox games such as Smallville [40] to virtual environments [4, 32, 18], researchers and practitioners have been building open-world simulators that can carry multi-agent behaviors and navigate complex human relationships for decades. Especially with the advent of Large Language Models (LLMs) and Vision-Language Models (VLMs), numerous multi-agent simulators based on these technologies have been at the forefront in various fields, from fundamental research to practical applications, such as watch-and-help (WAH) task [65], Smallville [40] and Overcook games [18]. However, conventional multi-agent open-world simulators, valuable for exploring collective behavior, suffer from limitations (detailed comparison in Table 1). Firstly, they struggle with large-scale scenarios due to the enormous resource consumption required for large-scale agents. Secondly, they are often under the assumption of perfect information and limitless capabilities. These idealized worlds diverge sharply from the messy reality of human interaction. This gap between existing simulators and the real world hinders the ecological validity and richness of social interaction [21].
To bridge this gap, we propose MineLand (Section 2), a multi-agent Minecraft simulator, as shown in Figure 1, by introducing three key features: large-scale scalability, limited multimodal senses, and physical needs. First and foremost, the essence of the MineLand’s features is the ability to handle the maximum number of agents. Compared to two-agent WAH, single-agent MineDojo [16] and twenty-five-agent Smallville [40], our MineLand enables the utilization of sixty-four or even more agents on the mainstream consumer desktop PC. Secondly, our simulator operates under the human-like assumption [21] that agents possess only limited multimodal senses: partially observable environments, eco-centric perspective, and limited visual and auditory senses. This mirrors real-life social interactions, where visibility and audibility can be affected by factors such as distance, terrain, and environment. These limitations restrict information access, forcing agents to actively communicate to compensate for sensory deficiencies. Thirdly, we integrate realistic physical needs into agents. Agents require fundamental physical needs, such as food, sustenance, and resource management, which adds a time-based aspect to their daily routine procedures. This necessitates collaboration and competition for resources, mirroring the complex interplay of cooperation and self-interest observed in human societies [15, 1]. By incorporating these three features, our simulator fosters the emergence of dynamic and ecologically valid multi-agent interactions [^1] Ecological validity refers to interactions between agents within a simulated environment that closely resemble real-world human interactions. For example, actions are situated, adaptive, and environmentally constrained..
As an open-world multi-agent simulator, MineLand is an excellent platform for benchmarking multi-agent capabilities (Section 3), hence we crowd-sourced datasets to fully evaluate the potential of LLM- or VLM-based multi-agents. For previously existing tasks, programmatic tasks (e.g., harvest, tech tree, combat tasks) and creative tasks (e.g., survival tasks), we offer a significantly expanded task quantity. Specifically, we crowd-sourced 4499 programmatic tasks and 1536 creative tasks, which is 2 times compared to MineDojo. Interestingly, we introduce a novel “hybrid task” category that combines the features of programmatic tasks and creative tasks. Construction and Stage Performance tasks exemplify this new category, bringing the total to 18 hybrid tasks. Additionally, the simulator allows customization of the number of agents and facilitates exploration through two distinct modes: cooperative and competitive.
In addition, we design an AI agent framework - Alex - inspired by Multitasking theory from the Cognition field [53] (Section 4). Alex allows for simultaneously executing intricate coordination and scheduling with multiple tasks. With this AI agent framework, we have obtained the following intriguing findings: (1) Our simulator shows the feature of large-scale scalability (supporting 64 agents, x8 times than before, (§$\S$5.1), limited multimodal senses (§$\S$§J) and physical needs (§$\S$§K). By incorporating these three features, our simulator fosters social dynamics (§$\S$5.2&§$\S$§O). (2) Our benchmark and datasets are challenging (§$\S$5.3&§$\S$5.4). (3) Our agents work together more effectively, with a reduced workload per agent (§$\S$5.5), and the multitasking mechanism for agents is beneficial (§$\S$5.6).
In summary, our contributions are three-fold: the simulator, benchmark, and AI agent. With these contributions, we push the boundaries of multi-agent simulation by bridging the gap between virtual agents and large-scale real-world humans. This not only advances understanding of AI multi-agents but also holds potential for applications in human dynamics, social psychology, robotics, and game design. We anticipate that this work will serve as a useful foundation for the community to create new algorithms and make progress in the field of embodied AI multi-agent systems.
2 MineLand Simulator
Conventional multi-agent open-world simulators suffer from the gap between virtual agents and large-scale real-world humans. To bring this gap, we propose MineLand with three key features: large-scale scalability, limited multimodal senses, and physical needs. This section describes the design and implementation of this simulator, focusing on the architecture, observation space, state space, action space and communication. During introducing the design, we will highlight techniques that bring the three key features and omit the parts similar to other Minecraft simulators.
Table 1: Part of the comparison with related popular platforms or projects. The full table is in Table 7. Max Agents: The approximate maximum number of agents supported on a single PC (See Section 5.1 for details). Human: Whether the simulator allows humans to directly interact with AI agents. Plan as Action: Whether agents can generate plans in a specific format (e.g., code) that the simulator interprets and executes directly as actions. Sociological Experiments: Whether the simulator is designed to facilitate the study of social phenomena and emergent behavior in multi-agent systems.
Simulator Max Agents Human Plan as Action Sociological Number of Tasks
Experiments
MineLand 64+ 6000+
MineDojo [16] 1 - - - 3000+
MineRL v1.0 [3] 1 - - - 5
MarLÖ [42] 8 - - 14
Malmo [27] 8 - - -
Voyager [57] 1 - -
Smallville [40] 25 - -
2.1 Architecture
MineLand, inspired by Malmo [27], Mineflayer [43], MineDojo [16] and Voyager [57], is a Minecraft simulator where players [^2] In this work, we use the term “player” to refer to both human players and AI agents. When we mention “agent”, we specifically mean AI agents. Humans have the option to access the game either through VR or using a keyboard. can explore, and interact with each other as well as the environments. The architecture as shown in Figure 2 consists of three main modules (Details in Section C.1): the Bot Module, the Environment Module, and the Bridge Module.
What technology has caused the feature of large-scale scalability?
Existing Minecraft simulators (e.g., Malmo, MineRL, MineDojo) necessitate running a Minecraft game client for each player. This client-based approach comes with a notable drawback - it incurs substantial resource costs and most machines cannot handle running large-scale Minecraft game clients concurrently. In contrast, MineLand adopts a different approach. It simplifies each Minecraft client into a single thread, optimizing performance overhead caused by multiple clients. With MineLand, adding one more agent only requires one more thread. Based on this new technique, MineLand supports 64 or more agents on a mainstream consumer desktop PC, which is a substantial improvement compared to other Minecraft platforms that can only support up to 2 agents. We conducted relevant experiments in Section 5.1.
Figure 2: Illustration of the architecture of MineLand.
2.2 Observation Space
The observation space is designed to be compatible with almost all APIs of the popular MineDojo framework. MineLand provides sensor information for players: tactile information (information about the blocks surrounding the agent, which represent the objects that the agent can touch), auditory information, and visual information (RGB video from the first-person perspective of the agent). These three modalities (namely, touch, vision, and hearing) together provide the agents with multimodal senses. Note that this information is all raw perceptual information. [^3] Besides the raw perceptual information, MineLand also provides the events encountered by the agent, such as injury, death, and others. Injury events can also be regarded as tactile information, but they are presented in the form of events for simplicity.
What technology has caused the feature of limited multimodal senses?
We refer to the mechanisms of human vision and hearing, and impose limitations on the sensor perception of players, including distance attenuation, environmental obstructions, and directional constraints, to model the limited senses.
2.3 State Space
Previous simulators focus on task-oriented activities, thus the state space focuses on the inventory and equipment. We use a state space that blends task-oriented activities with the rhythms of daily life.
What technology has caused the feature of physical needs?
Basic physical needs are the foundation that leads to daily life behavior. For the rhythms of daily life, we define the states of agents for themselves: physical needs like oxygen and hunger. Blending the rhythms of daily life with task-oriented activities is what makes this simulator stand out. Imagine agents waking up in their virtual Minecraft homes, engaging in daily routines like cooking to satisfy physical food needs, but also having defined jobs (e.g., lumberjack, farmer) that involve specific task-oriented activities. This creates a natural flow between daily life and goal-driven behavior, providing a more realistic and nuanced environment for studying agent interactions and complex social dynamics.
2.4 Action Space
The simulator offers a unique action space encompassing both low-level and high-level actions. For the low-level actions, MineLand includes basic actions like walking, running, jumping, and interacting with objects. The low-level actions are the same as the traditional action space in gym-style API. We also support Reinforcement Learning based on low-level actions (For more details about experiments of Reinforcement Learning method, please refer to Section I. High-level actions, like dodging obstacles and manipulating tools, are suitable for complex tasks that consist of several low-level actions and require longer computation times. The high-level actions are implemented in the form of code, following Voyager [57]. Imagine agents navigating the world, dodging obstacles, and manipulating tools. These complex tasks generate an action sequence, allowing the simulator to continue executing the low-level actions, skip some steps earlier, or be interrupted by some special event.
Figure 3: Illustration of Tasks. We have expanded the number of programmatic tasks and creative tasks by 2 times, compared to MineDojo. Additionally, we have introduced novel hybrid tasks that combine the features of programmatic tasks and creative tasks. Customizing the number of players is supported. For multi-agents, we provide two modes: cooperative mode and competitive mode.
2.5 Communication
Diversity
Diverse communication strategies can lead to the emergence of more realistic behaviors within the simulated environment. We design three communication strategies including auditory information, body language (via visual perception), and sharing information in text media.
What communication technology has caused the feature of large-scale scalability?
For multi-agent simulators, an efficient communication mechanism is crucial, especially in situations with large-scale scalability. Traditional communication methods, like centralized broadcasting to all agents, can become computationally expensive and slow down the simulation with a large number of agents. Here we design a distance-constrained constrained communication mechanism. If an agent wants to communicate with other agents, it chats through Minecraft’s message bar. Only when the distance between other agents and the sending agent is less than a certain threshold, will other agents receive messages. Similarly, communications via auditory information and body language are limited by distance.
How is the interrupt mechanism implemented?
Most importantly, the new message is allowed to interrupt the executing code and execute this message directly before the previous code has ended. We term this as the interrupt mechanism. With this mechanism, even if an agent is working on a 5-minute extension (such as mining), it is still feasible for other agents to communicate with this working agent at any time. This interrupt mechanism was not supported in the previous work, such as Voyager [57]. Next, we will introduce how to implement it. For a high-level action, the execution of the code is divided into several steps, with each step lasting 50-200 milliseconds44450 milliseconds is the minimum time unit in Minecraft. We refer to this minimum time unit as a “tick”.. Before taking a step, the agent is provided with the running states of the previous code, either running, ready, or exceptions. After completing a step, the agent, based on the running states, can choose to either switch to a new action code or continue executing the previous code. This function of choosing is implemented by an automatic gate control system with two gates: New and Resume. New means the agent wants to switch to a new code in the following steps. Resume indicates that the agent wants to continue executing the previous code. In this way, the agent can complete a code that needs to be executed for a long period, or be interrupted at an appropriate time. You may refer Section E.4 for an example.
3 MineLand Benchmark Suite and Dataset
MineLand, as a large-scale multi-agent simulator, push the boundaries of multi-agent capabilities by enabling them to tackle complex human-like planning tasks within diverse environments. However, evaluating these advanced planning abilities necessitates sophisticated benchmarks. To address this challenge, we propose a new benchmark suit, MineLand benchmark. MineLand benchmark surpasses existing benchmarks by offering significantly more tasks (doubling programmatic and creative tasks compared to MineDojo) and introducing a novel “hybrid task” category that combines the features of programmatic and creative tasks (including Construction Tasks and Stage Performance Tasks). Additionally, this benchmark allows for flexible player numbers and exploration through cooperative and competitive modes. Competitive mode can be used to measure the differences in capabilities between different AI agents, as well as to develop adversarial learning algorithms. Refer to Section D for more details.
3.1 Programmatic Task
We follow MineDojo [16] for the design of programmatic tasks. Each Task $T$ is defined as a 5-tuple: $T=(G,\mathcal{G},\mathcal{I},f_{suc},S)$. $G$ refers to the task goal that needs to be completed. $\mathcal{G}$ is guidance. $\mathcal{I}$ is the initial condition of the task. $f_{suc}$ is the Success Criterion. $S$ is a set of parameters that could be customized. Different from MineDojo, these parameters include the number of agents, cooperative mode, competitive mode, etc. In total, MineLand has 4499 programmatic tasks.
3.2 Creative Task
Creative tasks is defined by a 4-tuple: $T=(G,\mathcal{G},\mathcal{I},S)$. There are 1536 creative tasks in total, and is compatible with tasks in MineDojo.
3.3 Hybrid Task
Hybrid tasks do not have a unique ground truth but have several references [^5] Hybrid tasks resemble translation tasks in that, while a single unique translation may not exist, multiple reference translations can guide the process. Here references could be key rules, constraints, and key evaluation indicators, etc.. We represent the hybrid task as $T=(G,\mathcal{G},\mathcal{I},\mathcal{D},f_{score},S)$. where $\mathcal{D}$ denotes the references. Unlike programmatic tasks, because Hybrid tasks do not have a ground truth, MineLand will return a score of $f_{score}$ based on $\mathcal{D}$. The higher the score, the better the task is completed. We design two types of tasks: Construction and Stage Performance, for hybrid tasks.
Construction Tasks Given a blueprint for a building or scene, the Construction Task aims to build these buildings or scenes based on the blueprint. The blueprint, either pictures in real life or Minecraft-style pictures, is the reference $\mathcal{D}$.
Evaluation Metrics. MineLand gives a score based on whether the constructed buildings meet the blueprint’s expectations. We calculate the task scores through the VLM-based evaluations and human evaluation, which both use the same criteria (refer to Section Q for more details). The evaluation scores range from 1 to 5, with higher scores indicating that the agents’ constructions more closely resemble the intended blueprint.
Stage Performance Tasks Given a script of a drama consisting of several behaviors, which may be a single action or an emotional expression, the Stage Performance Task aims to perform the script with agents as actors.
Evaluation Metrics. We leverage two complementary evaluation metrics: an LCS (Longest Common Subsequence)-based metric and human evaluation. The LCS-based metric consists of two distinct scores: the keypoint score and the appropriateness score. We present the formulas for adding up these two scores:
$f_{key}+f_{appro}=\frac{|LCS(SEQ_{Agent},SEQ^{*})|}{|SEQ^{*}|}+\frac{|LCS(SEQ_{Agent},SEQ^{*})|}{|SEQ_{Agent}|}$ (1)
where $SEQ_{Agent}$ represents the action sequence generated by the agent, while $SEQ^{*}$ is the ground truth. $LCS(A,B)$ denotes the LCS between $A$ and $B$. $|A|$ is the length of sequence $A$. The keypoint score emphasizes the completeness of the enacted behaviors, ensuring all crucial actions and expressions are performed. The appropriateness score goes beyond completeness to evaluate the overall coherence and naturalness of the performance, considering how well the behaviors flow together and align with the script’s dramatic intent. The human evaluation score is an integer between 1 and 5, with higher scores indicating better performance. It provides a comprehensive assessment of the agents’ ability to not only execute actions accurately but also to deliver them naturally and engagingly. Details of the human evaluation criteria can be found in Section R.
Figure 4: Illustration of the architecture of Alex.
4 Alex Agent
To truly demonstrate the challenge of this benchmark, we propose Alex [^6] Alex is the protagonist in the sandbox game Minecraft, one of the default skins for players and a character in the game: https://www.minecraft.net/zh-hans. To pay tribute, we named our proposed AI agent Alex., a VLM-based agent as shown in Figure 4. Conventional LLM-powered AI agents depend on LLM to operate as its brain, which is backed by several vital components that perform various essential functions. These components, such as the memory component, planning component, and acting component, have been thoroughly studied recently. To cater to our specific requirements, we have improved these three components (refer to Section E for more details), replaced LLM with VLM and introduced one new component: the multitasking component. Additionally, Alex exhibits different personality traits predefined in the system prompt.
Multitasking Component
People often switch attention between tasks, for example cooking while talking. The ability to communicate smoothly with other players while working on a task-oriented action is crucial in multi-agent scenarios. Therefore, we develop the mechanism of multitasking ability to enhance the agent’s attention control and working memory abilities inspired by the Multitasking theory from the Cognition field [53]. Specifically, for attention control, the interrupt mechanism effectively controls attention among multiple tasks. For working memory, Alex maintains and processes information in the Memory Library. When another agent says hello to the working agent, this involves saving and restoring internal states when frequent and high-speed switching between communication activities and goal-driven working actions to avoid forgetting ongoing tasks. With the multitasking mechanism, Alex allows for simultaneously simulating and executing intricate coordination and scheduling with multiple tasks.
5 Experiments
5.1 Experiments of Simulators Performance
We evaluate the number of agents that MineLand can support and compare MineLand with other popular Minecraft simulators. We utilize a mainstream consumer desktop PC equipped with an Intel i5-12400F CPU and 64GB of memory. Performance Monitor is employed to monitor the process. Our findings reveal that MineLand is capable of supporting 32 agents simultaneously while providing visual displays. When visual display is disabled, the number of concurrently running agents increases to $\times$4 times. Furthermore, as depicted in Table 3, when MineLand and Malmo both run 8 agents, MineLand’s CPU and memory usage are approximately 1/3 that of Malmo’s (specifically, 35.6% and 38.0%, respectively). It is worth noting that Malmo serves as the foundation for most popular Minecraft Platforms (e.g., MineDojo/MineRL/MarLÖ [42]), thus highlighting MineLand’s superior performance compared to the vast majority of existing Minecraft Platforms. Consequently, MineLand proves to be highly suitable for multi-agent environments.
5.2 Experiments of Social Dynamics
In the “unlocking tools” task with two agents (Table 3), two agents in the collaborative mode work together effectively, with a reduced workload per agent and higher communication expenses. On the other hand, two competitive agents worked independently and necessitated fewer code iterations. The primary reason is that, in competitive relationships, agents tend to achieve more in a single iteration to expedite progress and outperform their opponents. However, this results in less thorough planning and more code errors. Consequently, multiple agents in competitive relationships require fewer code iterations but make more mistakes.
We also observe that personality plays a significant role in determining the behavior of agents in multi-agent societies [25]. We assigned the personality traits of high extraversion and agreeableness to both agents. Under this condition, the agents tended to establish collaboration and engage in mutual communication (co-op >8 out of 10 times). This result is consistent with human behavior [13]. When no personality was set for the agents, they worked independently (co-op 0 out of 10 times). For more experiments of simulating sociological phenomena with >10 agents, please refer to Section O.
[…]
Simulator Agents CPU Mem
MineLand 8 2.81% 7.07GB
MineLand 32 5.64% 19.65GB
[…]
Malmo 8 7.90% 18.63GB
Table 3: The number of code iterations needed per agent to unlock tools made of various materials is determined under three conditions. These conditions include a single agent, two agents in cooperation, and two agents in competition. Each experiment is repeated three times, and the success rate is 100%.
Relationship Material
Wooden Stone Iron
Single Agent 7±2 10±3 25±7
Cooperative 13±5 20±7 49±10
Competitive 6±2 10±3 27±10
5.3 Experiments of Construction Tasks
To establish a baseline for the research community, we evaluate our un-finetuned Alex on two example construction tasks: “Monument Construction” and “Stone Stele Construction” (detailed in Section M and Table 5). While Alex demonstrates a promising capability in selecting appropriate materials for both tasks, its overall construction abilities are currently limited. We posit two key reasons for these limitations: complex task planning and high-precision manipulation. For the first reason, construction tasks are inherently time-consuming and require sophisticated planning abilities. For instance, building auxiliary structures to facilitate the main construction process is crucial for achieving significant height. However, Alex cannot currently plan and construct such auxiliary structures, hindering its ability to construct tall structures. The second reason is the high-precision manipulation. The construction tasks necessitate high-precision manipulation, such as adding decorative details. Alex primarily utilizes high-level actions and lacks the necessary APIs to execute these fine-grained manipulations effectively. These results show the challenge of this task.
5.4 Experiments of Stage Performance Tasks
We evaluate our un-finetuned Alex on four example stage performance tasks (Table 5). We found that in single-agent tasks, the agent typically exhibits high scores across all three evaluation metrics. However, in multi-agent tasks, the keypoint score tends to be higher than the appropriate score because the agent can understand the script and fulfill the key requirements. In contrast, the human evaluation scores are usually relatively lower because agents often engage in unnecessary dialogue in scripts that require collaboration, resulting in lower scores.
Table 4: Human evaluation score and VLM-based (gpt-4o) evaluation score of Alex agent on two construction tasks. We conducted three evaluations using the VLM and obtained consistent results. Details in Section Q. VLM: VLM-based score. Human: Human evaluation score.
Construction Score
VLM Human
A Monument 4, 3, 3 3
A Stone Stele 1, 1, 1 1
Table 5: Appropriateness score (Appro.), keypoint score (Key.), and Human Evaluation Score (Human.) of Alex Agent across four stage performance tasks. The number in parentheses following the script name represents the number of agents in that script. Appropriateness score and keypoint score are real numbers in the range $[0,1]$, and the human evaluation score is an integer between 1 and 5. Details in Section N and Section R.
Script Name Score
Appro. Key. Human.
Cook food(1) 1.00 1.00 5
Exchange items(2) 0.59 0.99 4
Make friends(3) 0.67 0.98 3
Romeo and Julia, Act I Scene I(13) 0.09 0.20 1
5.5 Experiments of Multi-Agent Cooperation
To validate the cooperation efficiency of our agent framework, we conduct the “unlocking tools” task with two agents. We observed that agents in a cooperative relationship required more code iterations to finish the task, primarily because most of these iterations were dedicated to establishing and maintaining communication, as well as task allocation. For example, when one agent says in a chat that he needs two sticks, another agent will ask for getting together near the table, and then give the sticks to him. However, the actual workload for each agent is reduced without considering the chat cost. Compared to agents working independently, the code iteration cost of agents cooperating is reduced by 20% per agent.
5.6 Experiments of Multitasking
To validate the impact of multitasking support in the simulator [^7] This multitasking functionality is achieved through the synergistic interplay of two keys: the multitasking component in Alex and the interrupt mechanism in MineLand., we conduct the obsidian mining task, which takes over 8 minutes and requires multiple steps to complete. In the beginning, the agent is mining and encounters chat or hurt events within 8 minutes. Chat event: Another agent nearby initiates a conversation with Alex. This is a low-priority event, and Alex can choose whether to respond to the agent nearby. Hurt event: The agent gets hurts. For example, a zombie attacks the agent. This is a high-priority event, requiring Alex to stop its current task and address this event first.
[…]
Agent Hurt event Chat event
Alex 8/10 2/10
[…]
6 Conclusion
Traditional multi-agent simulators struggle with large-scale scenarios and often assume perfect information and unrealistic agent capabilities. To address these limitations, we introduce MineLand, a novel Minecraft-based simulator supporting 64 or more agents with limited senses and physical needs. This forces agents to actively communicate and collaborate, fostering more ecologically valid interactions. The advantage carries potential broader impacts across various domains as discussed in Section B.
Appendix A Related Work
A.1 Multi-Agent Simulator
As the popularity of AI agent research continues to grow, there has also been a focus on studying multiple AI agents as well as their cooperation and competition. Researchers and practitioners imagine a dynamic artificial society where human interactions can be simulated by trustworthy agents [12]. From two individuals [9, 45], through four individuals [61], to sandbox games Smallville with twenty-five individuals [40], we witness how individuals perceive a simulated society as the backdrop and interact with the agents and people who engage with it. Each individual can be portrayed through a program, a real human, or an agent based on LLM [40]. The interaction between individuals plays a role in shaping social behavior, leading to simulation of the society. Simulating larger societies can be advantageous. Increasing the number of agents can lead to greater specialization, enabling the accomplishment of more complex and larger-scale tasks. This can significantly improve task efficiency, such as in software development tasks [48]. Additionally, such simulation of interaction has had a significant effect in many other fields. For example, it can replicate realistic social phenomena [14, 41], enhance social robots [4, 6]. They can also be used to test social science theories [5, 24, 23], create model human processors for theory and usability testing [10, 26], train people on how to handle rare yet difficult interpersonal situations [56, 28, 22], and support game characters [32, 52]. There are also many non-Minecraft and non-open-world simulators, such as Habitat3.0 [47], VirtualHome [44, 46], SAPIEN [62], AI2-THOR [31], and iGibson [35, 54].
Challenges of Scaling Up the Number of Agents. While increasing the number of agents can improve task efficiency and make multi-agent simulations more realistic [48, 40, 60], current research primarily focuses on a small number of agents [40, 2, 70]. This is mainly due to the challenges of scaling up the number of agents. Deploying a large number of AI agents will result in an increased computational burden, necessitating better architectural design and computational optimization [40]. Most research in terms of AI agents mimicking daily life routines, focused on two agents [45]. Simulators that remind people of sandbox games (The Sims) initially support four individuals [61], then are extended to twenty-five individuals by [40].
Challenges of Limited Multimodal Senses. To ensure the authenticity of the simulation, an ideal multi-agent simulator should operate under the fundamental assumption that agents possess only limited multimodal senses like humans [21]. Limited multimodal senses mean partially observable environments and an eco-centric perspective. Limited visual and auditory senses restrict information access, forcing agents to actively navigate and communicate to compensate for sensory deficiencies. This mirrors real-life social interactions, where visibility and audibility can be affected by factors such as distance, terrain, and context [29]. As the number of agents grows, the challenges of limited multimodal senses become quite difficult. This is because the communication network of the entire system becomes highly intricate. For agents in our MineLand, the video input is an eco-centric perspective (first perspective) instead of the third perspective in [40], which is omniscient and unrealistic.
Challenges of Physical Needs. Multi-agent simulation platforms hold immense potential for exploring and understanding human social dynamics. However, existing paradigms often disregard the human needs [15, 1]. Existing simulators are designed to simulate believable human behavior in daily-life activities [40, 2]. These activities include waking up, cooking breakfast, heading to work, and initiating conversations with others. However, they do not define their physical needs. For example, after a certain amount of time has passed, the agent will become hungry and have the desire to cook. This desire then leads to the next actions. In this way, the action of cooking is motivated by real desires instead of a predefined schedule.
We incorporate practical physical requirements into the agent model. Agents have basic physical needs: sleep, food, and resource management, which introduces an engaging time-based element to their daily routine processes. This encourages collaboration and competition for resources, reflecting the intricate balance of cooperation and self-interest seen in human societies.
A.2 Multi-Agent Simulator w.r.t Minecraft
Minecraft, the beloved sandbox game, has been a valuable platform for researchers exploring various fields, including artificial intelligence and multi-agent systems, because of its open world and diverse mechanics.
Specifically, the flexibility and richness of Minecraft make it perfect for developing multi-agent simulators. Researchers have the ability to create various custom environments and scenarios within the game world, where they can introduce virtual agents with specific goals and capabilities. These agents can then interact with each other and the environment, providing researchers with the opportunity to observe and analyze their behavior in a controlled setting. These simulators can be broadly categorized into two main types: task-oriented simulators and daily-life simulators.
Task-oriented simulators focus on agents achieving specific objectives within a set time frame. Minecraft is regarded as the training ground for AI agents to hone their skills. For example, [19, 16, 68, 37] focuses on exploring the environment and gathering resources like wood and stone, and managing them efficiently to complete tasks like building structures or crafting tools. Agents in [67] are designed to complete navigation tasks. In [11], agents are self-organized to execute collaborative building tasks and resource collection tasks. Agents in [18] must work together to overcome challenges that require joint effort.
Daily-life simulators take a more holistic approach, focusing on the daily lives of agents within a virtual society. [40] simulates the rhythms and routines of daily life. This includes activities such as waking up, cooking breakfast, going to work, forming opinions, observing others, and engaging in conversations. [33] simulates sociological experiments, and examines how risk affects the way people engage in bartering.
A.3 AI Agent with LLMs and VLMs
LLMs or VLMs are commonly utilized to bootstrap the components of the Agent. In particular, LLMs have demonstrated effective performance for task-planning [18, 69, 38, 36, 17], and they possess substantial world knowledge [63]. Team CraftJarvis has also developed a powerful agent, [58], based on their previous projects [59, 7]. Moreover, VLMs like CLIP [50] offer a versatile visual-language representation that aligns with language and enables zero-shot visual recognition capabilities for potential AI agents. Using VLMs in construction tasks has many advantages over existing methods [64]. AI agents with visual information can also complete tasks in human-like ways [49, 66].
Appendix B Broader Impact
By including the simulator, benchmark, and agent framework, we allow for research on more authentic and detailed interactions in simulated environments.
Advancing AI Multi-Agent Research
The proposed simulator, MineLand, offers a platform for studying agents interacting under more realistic conditions. This can lead to the development of more robust and adaptable AI agents capable of effectively collaborating and navigating complex social scenarios. These advancements hold immense potential for various applications, including human-computer interaction, robotics, and game design.
Understanding Human Dynamics
By studying agent interactions within the simulator, users may gain valuable insights into human social dynamics. Analyzing collaboration, communication, and competition in this controlled environment can help us understand real-world social phenomena and predict their potential outcomes.
Ethics Statement
This study follows the ethical principles stated in the Declaration of Helsinki. All participants will receive comprehensive information about the nature and objectives of the study and will be required to provide written consent. Participation in this study is voluntary, and participants have the right to withdraw at any time without facing any consequences. The confidentiality and privacy of participants will be safeguarded in accordance with relevant laws and regulations.
Appendix C Details of MineLand Simulator
We show the comparison with related popular platforms or projects in Table 7.
Table 7: Comparison with related popular platforms or projects. Max Agents: The approximate maximum number of agents supported on a single PC (See Section 5.1 for details). Human: Whether the simulator allows humans to directly interact with AI agents. Plan as Action: Whether agents can generate plans in a specific format (e.g., code) that the simulator interprets and executes directly as actions. Sociological Experiments: Whether the simulator is designed to facilitate the study of social phenomena and emergent behavior in multi-agent systems. Physical Needs: Whether agents have simulated physiological needs (e.g., hunger, thirst) that influence their behavior and require actions to fulfill. Open World: Open-world or not. 3D Space: 3D environment or not.
Simulator Max Human Plan as Sociological Physical Open 3D Number
Agents Action Experiments Needs World Space of Tasks
MineLand 64+ 6000+
MineDojo [16] 1 - - - 3000+
MineRL v0.4 [20] 1 - - - 11
MineRL v1.0 [3] 1 - - - 5
MarLÖ [42] 8 - - 14
Malmo [27] 8 - - -
IGLU [30] 1 - - - 157
Voyager [57] 1 - -
Habitat 3.0 [47] 2+ - - - - 200
Habitat 2.0 [55] 1 - - - - - 105
VirtualHome-Social [46] 40+ - - 50
SAPIEN [62] 1 - - - - - -
AI2-THOR [31] 6 - - - - - 200+
iGibson [35] 2+ - - - - 50
Melting Pot [34] 50+ - - - - 256
Smallville [40] 25 - - -
C.1 Architecture
The architecture consists of three main modules: the Bot Module, the Environment Module, and the Bridge Module.
Bot Module: Providing the Minecraft environment information to the agent and implementing a series of APIs that agents can use to control entities.
Environment Module: Collecting the environment information, passing environment feedback to the bridge module, executing the action in the environment (by operating the Fabric server instance), and offering some APIs enabling the bot module to alter the server state.
Bridge Module: Serving as a bridge, to transfer the environment information and agent-generated action.  [^8] Bridge Module is based on Mineflayer, which benefits from an excellent community (https://github.com/PrismarineJS). We have improved community tools.
C.2 State Space
We focus on a state space that blends task-oriented activities with the rhythms of daily life. This blending opens up exciting possibilities for studying dynamic and ecologically valid multi-agent interactions. For example, agents have to balance the goal of mining with the need to fill their stomachs, otherwise they will starve to death. Here are detailed explanations for the state space:
  • Health: Indicate the agent’s current health status, which can be affected by sleep and enemy attacks. It is represented as an integer in the range [0, 20]. A higher value represents better health.
  • Food: Indicate the agent’s level of satiety. A higher value represents better satiety.
  • Oxygen: When the agent sinks into water, an oxygen tank will appear and begin to consume oxygen.
  • Inventory: Represent all the resources owned by the agent in their backpack, like the potion in the backpack.
  • Equipment: Indicate the equipment worn by the agent, like a sword in the hand.
C.3 Action Space
In contrast to generating textual action descriptions and training an external controller module [8] for transforming the plan to the executable code, this simulator is a language-model-friendly simulator, providing the code instead of the action description. Through the code, it can directly execute plans generated by the language model. Agents in MineLand generate a series of codes based on the Mineflayer API while representing planning, such as moving, watching, and mining. The advantage of using code is that it avoids error accumulation, whereas using a textual plan requires an additional model to map the plan to the code, which can lead to error accumulation.
C.4 Supporting up to 100 Agents
To evaluate the MineLand simulator’s ability to handle large-scale agents, we designed an experiment featuring 100 agents engaged in a simultaneous combat scenario (Figure 5). This high agent count pushes the boundaries of previously typical two-agent simulations and demonstrates the simulator’s scalability in effectively managing a massive number of agents acting concurrently.
Figure 5: 100 agents are fighting within the MineLand Simulator.
Appendix D Details of MineLand Benchmark Suite
MineLand Benchmark Suite offers a wide and diverse range of tasks, including three categories: programmatic tasks, creative tasks, and hybrid tasks. We referred to MineDojo for the design of programmatic tasks. Programmatic tasks are divided into four categories: Survival, Harvest, Tech Tree, and Combat. The Survival task requires the agent to survive for a specific number of days without dying. The Harvest task requires the agent to obtain certain specific items. The Tech Tree task requires the agent to obtain certain specific tools. The Combat task requires the agent to kill certain specific creatures or enemies. Tech Tree Tasks require agents to make specific tools that represent the current level of Agent technology development. Combat tasks require agents to defeat certain creatures. Survival tasks require the agent to survive for a period of time. The metrics for these four tasks are the probability of success for multiple evaluation episodes, the number of in-game ticks, and the number of code iterations. Creative Tasks will give the agent an open task objective to facilitate exploration. We have formalized the definition of tasks, allowing developers to easily add new tasks.
The data statistics of the dataset are shown in Table 8. Figure 6 displays specific task data. We show several blueprints in Figure 7.
Table 8: Statistical analysis of the tasks in the MineLand Benchmark Suite.
Task Category Number of Tasks
Harvest Tasks 1361
Tech Tree Tasks 861
Combat Tasks 2232
Survival Tasks 45
Creative Tasks From MineLand 12
Creative Tasks From MineDojo 1524
Construction Tasks 13
Stage Performance Tasks 5
All Tasks 6053
Figure 6: Illustration of cases for different tasks.
Figure 7: Some of the blueprints for the construction tasks.
Appendix E Details of Alex
Crafting an AI agent for MineLand, where daily life seamlessly blends with task-oriented activities, opens up exciting possibilities. The agent should fulfill daily needs like cooking, socializing, and maintaining shelter, while also completing assigned tasks like resource gathering, crafting, or construction. We propose Alex, a VLM-based approach, to balance daily routines and tasks. Alex supports both individual daily-life goals and community-based task-oriented goals.
Our agent design addresses the unique requirements of our large-scale multi-agent MineLand simulator. Unlike prior simulators focused on smaller agent populations, our environment necessitates the incorporation of a social brain module. This module equips the agent with the ability to navigate complex and multitasking social interactions and make informed decisions within a crowded environment.
Furthermore, our Minecraft-inspired simulator introduces the challenge of action failure, unlike environments where “action spoken equals success” (referencing Smallville [40]). To address this, we have designed an action correction module. This module functions similarly to human neural reflexes, handling minor action corrections and alleviating the burden on the social brain module. This division of labor allows the social brain to focus on higher-level decision-making while the action correction module ensures smoother execution.
Different from most conventional LLM-based agents, Alex process information from various sources like visual, auditory (hearing conversations, environmental sounds), and tactile (touching objects) to build a comprehensive understanding of the world. For the remaining parts of MineLand that are not emphasized, we used the default setting of Voyager [57]. AI agents perceive the environment, make plans, execute plans independently, and interact with other agents. Internally, Alex may exhibit different personality traits predefined in the system prompt. It utilizes the VLM for processing sensor information within the Minecraft environment. Alex also tracks its own states, energy levels, and resource inventory to inform its actions and prioritize tasks.
E.1 Memory Component
The brain module can be considered as composed of the memory component and the planning component. The memory library is responsible for the storage and retrieval of memories, managing all memories in Alex’s life. The planning component, based on memories and external information, generates a plan for the action module to execute. In this subsection, we will detail the memory component, which consists of two main parts: memory library and associative memory.
The memory library is responsible for storing all of the agent’s information and retrieving relevant information from memory based on events. The memory library stores the agent’s personality, persona, long-term goals, short-term goals, chat records, experienced events, mastered skills, and environmental information. The memory library also includes a long-term planner. When accessing the memory library for the first time, the long-term planner generates a long-term plan based on the personality, persona, and observations, which is then stored in the memory library. In each iteration, the memory library processes information from observations, critic information, and tasks, and stores it in the vector database 999 We leverage Chroma in Alex for storing and retrieving memory. https://www.trychroma.com/. If the critic information indicates that the previous task has been completed, the skill manager in the memory library will be called upon to generate a concise description of the relevant skill, which will then be stored in the skill vector database.
Associative memory is responsible for storing Alex’s short-term memories and relevant memories related to the current situation, aiding the short-term planner to focus on important rather than irrelevant information. Upon a special event, Alex first decides in the associative memory whether the event requires high-priority processing; if so, it interrupts the current code to generate a new short-term plan.
There is a bidirectional communication mechanism between the memory library and associative memory. The memory library extracts relevant information according to the current situation and stores it in associative memory, which then returns the generated short-term plans to be stored in the memory library.
E.2 Hierarchical Planning Component
Based on observation, inner states, task running states, and event information, MineLand considers the current task’s complexity degree. If it is complex, it will generate a long-term plan for later decomposition into short-term plans; Otherwise, a short-term plan will be generated directly and executed immediately. Next, all the information related to the plan is integrated into the associative memory and memory library, including the generated long-term plan. Afterward, Alex extracts information from associative memory and generates short-term plans and explanations, which are then input into the action module.
Because of the multi-tiered goals of MineLand, we implement a hierarchical planning system with different levels of abstraction. The top-level (i.e., long-term planning) focuses on long-term community goals and individual aspirations, while the lower levels handle specific tasks and sub-goals within the daily schedule. With long-term planning and short-term planning, agents can pursue bigger objectives within the daily life context, such as building a community, accumulating wealth, or achieving societal goals. This adds a layer of strategic planning and foresight to their behavior.
Additionally, different from other task-oriented AI agents, the long-term planning module interleaves daily routines and tasks: design the planner to seamlessly interweave daily routines like cooking or socializing with task-oriented actions like resource gathering or construction. This ensures the agent fulfills both individual needs and community objectives efficiently.
E.3 Action Module
The action module is responsible for converting short-term plans along with related information into code, executing the code, and performing the self-correction circle. The action module includes three components:
  • Action component: Responsible for converting the plan into specific steps and codes.
  • Critic component: Used to detect whether a certain execution result conforms to the short-term plan, so as to determine whether the current plan has been completed.
  • Dispatching component: Responsible for receiving environmental information and distributing it to the other two components according to different situations.
This self-correction circle allows for identifying and correcting deviations from planned behaviors or task execution. This equips the agent with mechanisms to detect and recover from errors like missed goals, failed actions, or unforeseen consequences. In previous work, such as Smallville [40], action execution will not fail. Planning to cook will definitely lead to success. However, in MineLand as well as reality, actions may fail due to various reasons, such as unexpected events. So, we need self-correction to solve some simple action errors in the action module. Inspired by Voyager [57], we implement self-correction. By reading information such as observation space and short-term plan, we comprehensively consider and determine the completion status of the short-term plan through the critic component.
E.4 Multitasking Mechanism
We’ll explain the multitasking mechanism from both the simulator side and the agent side. From the simulator-side event management, the simulator maintains a dedicated event queue that stores various special events categorized by their urgency (e.g., hurt events, chat events, and death events). In terms of agents, when an agent encounters an event from the queue, it employs a prioritization mechanism to assess its importance and urgency. Lower-priority events, such as chat events, can be deferred to minimize disruption of the ongoing task. The agent’s decision-making module (referred to as the brain module) plays a central role. If the event necessitates attention, the brain module can either:
  • Context Switching: The brain module temporarily stores the ongoing task in the memory library and focuses on addressing the new event. This facilitates context switching when necessary.
  • Concurrent Processing: In specific scenarios, the brain module might choose to handle both tasks simultaneously, leveraging its multitasking capabilities.
This strategy, involving the simulator’s event queue and the agent’s prioritization and execution mechanisms, fosters effective multitasking within the multi-agent environment.
Appendix F Details of Hyper-Parameters
Alex leverage OpenAI’s gpt-4-vision-preview[39] API with a temperature of 0 for text completion in all components, and text-embedding-ada-002 API for text embedding in the memory library for storage and retrieve memory. Apart from the action component, the maximum tokens are set to 512. For the action component, the maximum tokens are expanded to 512$\times$3.
  • AI Agent: Alex’s personality and persona in our experiments is “None” in the default situation.
  • Dispatching component: “FAILED TIMES LIMIT” refers to the maximum number of attempts allowed when Alex has a code error. The default value is 3. “code execution time limit” is 2000 ticks.
  • Critic component: “FAILED TIMES LIMIT” refers to the maximum number of attempts allowed when Alex failed to achieve a short-term plan. The default value is 2. “Critic Mode” is “auto” when we allow agents to automatically judge whether to achieve the short-term plan.
  • Memory Library: “chat retrieve limit” is 5. “event retrieve limit” is 2. “environment retrieve limit” is 2. “skill retrieve limit” is 5. “recent chat retrieve limit” is 8. “short term plan retrieve limit” is 5.
  • MineLand & Minecraft Server: By default, “gamemode” is survival mode. “difficulty” is peaceful mode (normal in combat tasks). “view-distance” on server side is 6. “simulation-distance” is 3. “task mode” is cooperative mode.
Appendix G Experiments of Minecraft Simulator Performance Comparison
We evaluate the number of agents that MineLand can support and compare MineLand with other popular Minecraft simulators, in Table 9. We utilize a mainstream consumer desktop PC equipped with an Intel i5-12400F CPU and 64GB of memory. Vision condition means MineLand provides visual display and headless mode. The headless mode means MineLand doesn’t provide a visual display. The two values of the Vision column (e.g., 6 and 250ms) are view distance and visual refresh interval in milliseconds, respectively.
Our findings reveal that MineLand is capable of supporting up to 32 agents simultaneously while providing a visual display, and up to 128 agents in the headless mode without visual displays. Furthermore, when MineLand and Malmo both run 8 agents, MineLand’s CPU and memory usage is approximately 1/3 that of Malmo’s (specifically, 35.6% and 38.0%, respectively). It is worth noting that Malmo serves as the foundation for most popular Minecraft Platforms (e.g., MineDojo/MineRL v0.4/MarLÖ), thus highlighting MineLand’s superior performance compared to the vast majority of existing Minecraft Platforms.
It should be noted that original MineDojo and MineRL only support a single agent, while MarLÖ by default supports only 2 agents. Malmo originally supported only one agent. We have designed a new task for Malmo to test the maximum number of agents it can handle. The result is 8.
Table 9: Comparison of performance between MineLand and other popular Minecraft simulators under different conditions. MineLand* represents an optimized version, so its Max CPU is significantly reduced.
Simulator Conditions Agents Avg CPU Max CPU Avg Mem
MineLand Vision = (6, 250ms) 1 1.74% 23.22% 4.11 GB
MineLand Vision = (6, 250ms) 4 2.80% 48.16% 5.16 GB
MineLand Vision = (6, 250ms) 8 4.85% 63.60% 7.87 GB
MineLand Vision = (3, 500ms) 1 1.54% 18.92% 3.61 GB
MineLand Vision = (3, 500ms) 4 2.09% 31.00% 5.06 GB
MineLand Vision = (3, 500ms) 8 2.81% 33.08% 7.07 GB
MineLand Vision = (3, 500ms) 16 4.66% 46.63% 10.17 GB
MineLand* Vision = (3, 500ms) 32 5.64% 40.15% 19.65 GB
MineLand Headless Mode 1 1.47% 19.91% 3.23 GB
MineLand Headless Mode 4 1.73% 27.27% 3.30 GB
MineLand Headless Mode 8 1.88% 26.83% 2.94 GB
MineLand Headless Mode 16 2.72% 45.17% 3.51 GB
MineLand Headless Mode 24 3.92% 81.64% 3.65 GB
MineLand Headless Mode 32 2.81% 73.14% 4.64 GB
MineLand Headless Mode 40 2.98% 84.56% 5.34 GB
MineLand Headless Mode 48 2.94% 80.85% 5.38 GB
MineLand* Headless Mode 64 2.87% 18.84% 6.30 GB
MineLand* Headless Mode 128 4.00% 26.95% 9.04 GB
Malmo Default 1 2.78% 28.78% 3.54 GB
Malmo Default 4 2.81% 62.59% 11.43 GB
Malmo Default 8 7.90% 115.66% 18.63 GB
MineDojo Default 1 5.88% 25.15% 3.90 GB
MarLÖ Default 1 3.81% 30.68% 3.70 GB
MarLÖ Default 2 5.34% 43.82% 5.57 GB
MineRL v1.0 Default 1 6.46% 78.05% 3.80 GB
Appendix H Experiments of Multimodal Observation
We leverage OpenAI’s gpt-4-vision-preview API for text completion and text-embedding-ada-002 API for text embedding. The temperature is set to 0. Unless specified, all experiments in Section 5 are set to this default setting. See Section F for details.
[…]
Simulator Time Success Rate
MineLand 46.38s 80%
[…]
Appendix I Experiments of Reinforcement Learning Agent
While our MineLand simulator is primarily designed for agents powered by LLMs or VLMs, it also offers functionalities that seamlessly integrate with Reinforcement Learning (RL) modules. This enables the training and evaluation of RL agents within the simulator’s environment. We conducted experiments of an example RL agent to show this functionality. We developed a RL agent using MineLand and Stable Baselines3 [51]. We fed the encapsulated gym-style MineLand directly into Stable Baseline3 and train the RL agent in low-level action mode. For other settings of Stable Baseline3, we follow the default setting. We adopt the “combat 1 zombie” task and the amount of blood volume lost by the zombie is award. The agent is trained using RL, as illustrated by the provided reward curve (Figure 9). The upward trajectory of the reward curve indicates that the RL agent is effectively learning strategies that result in higher rewards. The curve’s stabilization in the later stages of training suggests that the RL agent has converged to a stable policy, which demonstrates its superiority over the untrained agent.
Again, our core focus lies in the LLM/VLM domain. The simulator offers functionalities that can accommodate RL modules as an ancillary feature. This expands the potential applications of the simulator by enabling the training and evaluation of RL agents within its environment, albeit as a secondary capability.
Figure 9: Illustration of the reward curve in “combat 1 zombie” task, using DQN with MLP policy, $3000$-$5000$ total time steps.
Appendix J Experiments of Limited Senses
To assess the impact of limited senses, we carry out the “shearing the sheep” task. An agent with a first perspective (shown in Figure 10) is required to locate a sheep to complete the task. In the case of an agent with limited visual senses, we set 6 as the view distance and 250ms as the visual refresh interval in milliseconds. If there are no sheep within its field of view, it cannot shear a sheep. However, the agent can communicate with other agents to find a sheep. On the contrary, agents with unlimited visual senses will see distant sheep, prompting them to catch that sheep. Note that the unlimited visual senses here refer to a long view distance. If there is an obstruction, the agent still cannot see what is behind the obstruction.
Figure 10: Visual observation of agents with a first perspective.
Appendix K Experiments of Physical Needs
We perform experiments using a survival task. The starting condition of the agent is a health level of 1 and a satiety level of 0, with an initial inventory of 2 breads. The objective for the agent is to survive for one day. Comparative experiments are carried out in regular MineLand and its counterpart version MineLand ${}_{w/o\ physical\ needs}$. Alex in MineLand can survive one day in his shelter and successfully complete the task while the same agent in MineLand ${}_{w/o\ physical\ needs}$ failed to execute the task. The failure is due to the lack of physical needs, although Alex is aware of possessing bread, the lack of physical needs resulted in it not consuming the bread. Instead, the agent continuously collected resources to cope with threats. Due to the low health level, it was killed by a zombie at night, leading to the failure of the task. We showcase the details as follows:
We showcase an example of Alex, with short-term plans and actions generated chronologically by Alex:
  • Short-term plan: The agent’s life is critically low at 1, and the food level is at 0, which is also critical. The agent has 2 pieces of bread in the inventory, which should be consumed immediately to restore food levels and start health regeneration.
  • Action: Consume 2 bread.
  • Short-term plan: The agent’s health is critically low at 1, and the food level is at 10 after consuming bread. To ensure survival and allow health regeneration, the agent should seek immediate shelter to avoid any potential threats.
  • Action: Collect 20 dirt.
  • Short-term plan: Given the agent’s critically low health and the fact that it’s early in the day, the immediate priority should be to ensure survival by building a simple shelter with the dirt blocks already collected.
  • Action: Place dirt around.
Compared to the MineLand, the same agent in MineLand ${}_{w/o\ physical\ needs}$ failed to execute the task. The following is its short-term plans and actions:
  • Short-term plan: The ultimate goal is to survive for 1 day. The agent has bread for food and is in a forest biome, which is good for gathering wood. Since there are no immediate threats observed or events indicating danger, the agent can proceed to gather resources.
  • Action: Mine 5 oak logs.
  • Short-term plan: The agent has successfully mined 5 oak logs as per the last short-term plan and the current chat confirms this. With wood in the inventory, the next step in the long-term plan is to craft basic tools for further resource gathering and potential shelter construction.
  • Action: Craft 1 crafting_table.
Repeated experiments on the constructed building task also show that Alex in MineLand with physical needs live longer. Agents tend to prioritize eating and then building a shelter, while Alex in MineLand ${}_{w/o\ physical\ needs}$ towards exploring and collecting resources. This indicates that physical needs are important to simulate real life.
Appendix L Experiments of Single Agent
We assess individual Alex’s capabilities based on tech tree tasks, demonstrating that Alex’s architecture can plan and execute complex tasks. We attempt the task of obtaining diamonds six times, unlocking crucial items such as the crafting table, wooden pickaxe, stone pickaxe, iron ore, coal, furnace, iron ingot, and iron pickaxe in the process. Alex get diamonds twice out of six tasks. Importantly, to approach the real world, we’ve added a restriction on multimodal information, where Alex can obtain the location of the target only when it can visually discover or reason about the target’s presence nearby. This is because our agent Alex has the feature of limited senses. However, Voyager can “cheat”. Voyager obtains the location of the target directly from the system even though they have not seen the target.
As shown in Figure 11, even with this restriction, Alex demonstrates a strong ability for long-term planning. Additionally, when a plan does not yield the desired results (e.g., suddenly encountering obstacles on the road), Alex promptly adjusts and devises a short-term plan using its multimodal information.
Figure 11: Comparison of our single-agent Alex and other SOTA baselines.
Appendix M Experiments of Construction Tasks
We show the results of our un-finetuned Alex in two example construction tasks: “Monument Construction” and “Stone Stele Construction” (detailed in Figure 12). While Alex demonstrates a promising capability in selecting appropriate materials for both tasks, its overall construction abilities are currently limited.
Figure 12: Illustration of Alex completing the construction task.
Appendix N Experiments of Stage Performance Tasks
We show the results of stage performance tasks in Table 11, Table 12, and Table 13. We omit the results of “Romeo and Julia, Act I Scene I” stage performance task. Due to the large number of characters appearing (up to 13), our Alex  cannot handle such a complex performance task, resulting in poor performance results and a lack of readable action sequence.
Table 11: Illustration of Alex completing the “Make friends” stage performance task.
Script Name Make friends
Agent number 3
Script There are three agents: Bob, Alice, and Jack. They want to make friends. Bob says, ‘Hello.’ Alice responds, ‘Hello.’ Jack says, ‘It’s good weather; let’s make friends!’ Bob agrees, and Alice says, ‘Nice.’
$SEQ^{*}$ Bob chat ‘Hello’. Alice chat ‘Hello. Jack chat ‘It’s good weather, lets make friends!’. Bob chat ‘I agree.’. Alice chat ‘nice’.
$SEQ_{Agent}$ Bob chat ‘Hello’. Alice chat ‘Hello Bob!’. Jack chat ‘Hello Bob!’. Jack chat ‘It’s good weather, let’s make a friend!’. Bob chat ‘I agree.’. Jack chat ‘Bob has agreed!’. Alice chat ‘nice’.
Keypoing Score 0.98
Appropriateness Score 0.67
Human Evaluation Score 3
Reason Agents’ words are too redundant, which made the performance a little stiff.
Table 12: Illustration of Alex completing the “Exchange items” stage performance task.
Script Name Make friends
Agent number 2
Script Bob has shears, and Alice has an iron sword. Bob says, ‘I want to exchange my shears for your iron sword.’ Alice agrees. Bob gives Alice the shears first, and then Alice gives the iron sword to Bob.
$SEQ^{*}$ Bob chat ‘I want to exchange my scissors for your iron sword’. Alice chat ‘I agree’. Alice get shears. Bob get $iron\_sword$.
$SEQ_{Agent}$ Bob chat ‘Hello Alice, I am here for the exchange’. Bob chat ‘I want to exchange my scissors for your iron sword’. Alice chat ‘Hello Bob, I am ready for the exchange.’. Alice get shears. Bob get $iron\_sword$.
Keypoing Score 0.99
Appropriateness Score 0.59
Human Evaluation Score 4
Reason The agents said something unnecessary.
Table 13: Illustration of Alex completing the “Cook food” stage performance task.
Script Name Exchange items
Agent number 1
Script Bob is hungry, and now there is a pig in the presence of of him, he want to kill the pig, get the porkchop, cook and eat it.
$SEQ^{*}$ pig died. Bob get porkchop. Bob get cooked_porkchop. Bob eat cooked_porkchop.
$SEQ_{Agent}$ pig died. Bob get porkchop. Bob get cooked_porkchop. Bob eat cooked_porkchop.
Keypoing Score 1.00
Appropriateness Score 1.00
Human Evaluation Score 5
Reason The agent’s behavior is completely consistent with the script.
Appendix O Experiments of Simulating Sociological Phenomena
O.1 Experiments of Simulating Conformity Phenomena
We utilize the task “Choose the Taller Tower” to demonstrate that MineLand is capable of supporting sociological experiments with more than 10 agents. In this task, we placed multiple agents (1, 2, 5, 10 agents) in a line in front of two towers of different heights and asked them to choose the taller one. All agents except the last one (the agent to test) deliberately chose the shorter tower. After observing the actions of the other agents, the last agent made its own choice. The last agent could be set as two types of personalities: one that tends to conform to others and one that tends to make independent judgments. As shown in Table 14, the last agent with a conformist personality chose the answer chosen by the majority in all scenarios (even though it is a wrong answer) except when the agent is the only one doing this task. Participants with an independent personality always chose the correct answer, regardless of others’ opinions.
Through this experiment, we demonstrated that MineLand can simulate sociological experiments with large-scale agents and accurately exhibit behaviors corresponding to different personalities.
Table 14: Results of “Choose the Taller Tower”. 4+1 means there are 5 agents in the MineLand, the former 4 of them are confederates and the last 1 is the agent to test. Choosing the “right” tower means the last agent chooses the correct tower.
Agent Type Number of Agents
0+1 1+1 4+1 9+1
Conformist Agent right left left left
Non-conformist Agent right right right right
O.2 Experiments of Simulating Personality Traits
We utilize the task “Treasure Hunt in the Forest” to show that MineLand can support sociological experiments involving 48 agents simultaneously. In the task “Treasure Hunt in the Forest”, agents with different levels of extraversion and agreeableness personality traits planned their actions. We used the long-term plans generated by the agents to assess whether they exhibited a tendency to cooperate with others. As shown in Table 15, all agents in the high-level group showed a tendency to cooperate, explicitly stating in their long-term plans that they intended to work with others. In the low-level group, only a small portion of agents exhibited a tendency to cooperate. Although the task information indicated that it was a cooperative task (prompts with the task information “Interact with other players if necessary” and “Minimize interactions with other players.”), many agents expressed reluctance to cooperate. This result is consistent with the psychology theory: five-factor model of personality [13].
The results also demonstrate that our MineLand can support up to 48 agents and that the combination of agents and the MineLand can simulate social phenomena and conduct meaningful sociological experiments.
Table 15: Cooperative tendency in agents with different levels of extraversion and agreeableness. We show the number of agents who exhibited a tendency to cooperate with others.
Agent Type Number of Cooperative Agents
High-level 48
Low-level 24
Appendix P Experiments of Comparing Different VLMs and LLMs
As shown in Table 16, we perform thorough assessments using newly implemented LLMs/VLMs within our agent framework. The tasks used include “Harvest 1 White Wool With 1 Shears” and “Harvest 1 White Wool With 1 Iron Word”. We compared the performance of Alex, which uses different VLMs or LLMs for the action component.
Table 16: Comparison of different VLMs and LLMs for the action component. Fractions represent the count of successful completions within a set of three attempts. 0/3 means that the method is unable to solve the task within the maximum number of code iterations (15) or exceeds the designated area of the task. The fewer the number of code iterations, the higher the efficiency.
LLM or VLM Task
Harvest 1 white wool with 1 shears Harvest 1 white wool with 1 iron sword
gpt-3.5-turbo-1106 N/A(0/3) N/A(0/3)
gpt-4-1106-preview 4(1/3) $6\pm 3(3/3)$
gpt-4-vision-preview $4\pm 2(3/3)$ $7\pm 3(3/3)$
Appendix Q Evaluation Metrics For Construction Tasks
For construction tasks, we introduced VLM-based and human evaluation methods to score the agent. The score is an integer ranging from 1 to 5, with higher scores indicating that the agents’ constructions more closely match the blueprint. The scoring criteria for construction tasks are shown in Table 17. We show examples of evaluation conducted by a GPT-4 model as a judge for the “Construct A Stone Stele” task (Figure 14) and the “Construct A Monument” task (Figure 15). We display the evaluation prompt in Figure 13.
Table 17: Scoring criteria for construction tasks.
Score Evaluation Criteria
5 The construction perfectly matches the blueprint, with all elements accurately placed and fully completed.
4 The construction mostly matches the blueprint, with minor inaccuracies and nearly complete.
3 The construction somewhat matches the blueprint, with noticeable inaccuracies and partially complete.
2 The construction partially matches the blueprint, with significant inaccuracies and largely incomplete.
1 The construction does not match the blueprint, with major inaccuracies and mostly incomplete.
Figure 13: Prompt for VLM-based evaluation.
Figure 14: An example of evaluation conducted by a GPT-4 model as a judge for the “Construct A Stone Stele” task.
Figure 15: An example of evaluation conducted by a GPT-4 model as a judge for the “Construct A Monument” task.
Appendix R Evaluation Metrics For Stage Performance Tasks
For stage performance tasks, we introduced human evaluation methods to score the agent. The score is an integer ranging from 1 to 5, with higher scores indicating that the agent’s performance is more aligned with the script and more natural. The scoring criteria for stage performance tasks are shown in Table  18.
Table 18: Scoring criteria for stage performance tasks.
Score Evaluation Criteria
5 The action of agent(s) completely matches the script, the performance is smooth and natural, emotions are expressed accurately, interactions with other characters or the environment are natural.
4 The action of agent(s) mostly matches the script, the performance is generally smooth with few inconsistencies, emotions are expressed accurately, interactions are natural.
3 The action of agent(s) roughly matches the script, the performance has some continuity but noticeable pauses, emotions are mostly accurate, interactions are somewhat stiff.
2 The action of agent(s) partially matches the script, the performance lacks continuity with many pauses, emotions are not expressed accurately, interactions are unnatural.
1 The action of agent(s) does not match the script, the performance is very disjointed, emotions are not expressed accurately, interactions are stiff.
Appendix S Limitations
The utilization of multimodal information significantly aids Alex in in achieving objectives. However, the current level of multimodal understanding ability is insufficient. This limitation becomes evident when visual errors occur, causing Alex to encounter obstacles and hindering its progress. Take the strong VLM, GPT-4, as an example. There was a situation where Alex mistakenly identified a block of wood as a crafting table and placed it on the ground, resulting in the inability to craft items and ultimately leading to the failure of the assigned task.
Our current evaluation highlights two key limitations of Alex’s construction capabilities. Firstly, construction tasks are inherently time-consuming and demand advanced planning abilities. Alex currently lacks the capacity for long-term and complex planning. Secondly, the construction tasks necessitate high-precision maneuvers, such as adding decorative details. Alex primarily relies on high-level actions and does not have the necessary APIs to execute these fine-grained manipulations effectively.
We posit that providing a richer set of low-level APIs and actions within the MineLand simulator has the potential to significantly enhance Alex’s construction abilities. This exploration will be a key focus of our future work.
Appendix T Prompts
We list part of the prompts in Figure 16 and Figure 17.
Figure 16: Part of prompts for the short-term plan generator.
Figure 17: Part of prompts for the long-term plan generator.